OpenStudy (anonymous):

In 1992, the life expectancy of males in a certain country was 65.4 years. In 1998, it was 67.5 years. Let E represent the life expectancy in year t and let t represent the number of years since 1992. E(t)=_t+_ Use the function to predict the life expectancy of males in 2004.

OpenStudy (saifoo.khan):

(1992, 65.4) and (1998, 67.8) slope = 0.4 so E(t)= 0.4t+65.4

OpenStudy (saifoo.khan):

2004 - 1992 = 12 E(t)= 0.4t+65.4 E(t)= 0.4(12)+65.4 E(12) = 70.2